# Characterizing categoricity in the class $\mathrm{Add}(M)$

### v5

@@ Abstract
As an application, we show that the class of pure-projective $R$-modules is categorical in some (all) big cardinal if and only if $R$ is $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$; the class of semisimple $R$-modules is categorical in some (all) big cardinal if and only if $R$ admits a unique simple module up to isomorphism, partly answering a question proposed in [5, Mazari-Armida M., Characterizing categoricity in several classes of modules. J. Algebra 617, 382-401 (2023)].
Key Words: $\mathrm{Add}(M)$; pure-projecitve modules, categoricity, Shelah’s Categoricity Conjecture.
@@ 1. Introduction
[…]
@@
(((=Theorem 2.2))) Let $R$ be a ring and $M$ be an $R$-module which can be generated by $\leq\aleph$ elements. Then the following statements are equivalent:
@@
(((=Theorem 2.5))) Let $R$ be a ring, $S\SS$S the class of all semisimple $R$-modules and $S$ the direct sum of all simple modules non-isomorphic to each other. Then the following statements are equivalent:
@@
    $S\SS$S is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
@@
    $S\SS$S is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
@@
(((=Theorem 2.7))) Let $R$ be a ring and $P\mathcal{P}roj$ the class of all pure-projective $R$-modules. Then the following statements are equivalent:
@@
    $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$.
@@
Let $M$ be an $R$-module. Set $\mathrm{Add}(M)$ to be the class of all $R$-modules which is a direct summand of direct sums of copies of $M$. We call an $R$-module $N$ is $M$-free if $N\cong M^{(\kappa)}$ for some cardinal $\kappa.$ The following main theorem shows that the categoricity of $\mathrm{Add}{M}$ in a big cardinal is equivalent to its categoricity in a tail of cardinals.
@@
(3)(1)(2)(3$\Rightarrow$(1$\Rightarrow$(2)( 3 ) ⇒ ( 1 ) ⇒ ( 2 ) and (3)(4)(2)(3$\Rightarrow$(4$\Rightarrow$(2)( 3 ) ⇒ ( 4 ) ⇒ ( 2 ): Trivial.
(2)(3)(2$\Rightarrow$(3)( 2 ) ⇒ ( 3 ) Suppose $\mathrm{Add}(M)$ is $\kappa_{0}$-categorical in some fixed cardinal $\kappa_{0}>\|R\|+\aleph+\aleph_{0}$. We will show every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is $M$-free by dividing into two cases:
Case I: $\kappa\geq\kappa_{0}.$ Let $P$ be an $R$-module with cardinal $\kappa$ in $\mathrm{Add}(M)$. Then $P=\bigoplus\limits_{i\in I}P_{i}$ where each $P_{i}$ is $(\leq\kappa)$-generated $R$-module by Lemma 2.1. Since $\|P\|=\kappa>\|R\|+\aleph+\aleph_{0}$ and $\|P_{i}\|\leq\|R\|+\aleph+\aleph_{0}$ for each $i\in I$, we have $|I|=\kappa.$ Then we can give a partition $\{I_{\lambda}\mid\lambda\in\Lambda\}$ of $I$ with each $|I_{\lambda}|=\kappa_{0}$. Set $\bigoplus\limits_{i\in I_{\lambda}}P_{i}=P^{\prime}_{\lambda}$. Then $P=\bigoplus\limits_{\lambda\in\Lambda}P^{\prime}_{\lambda}$. Note that $\|P^{\prime}_{\lambda}\|=\kappa_{0}$. So $P^{\prime}_{\lambda}\cong M^{(\kappa_{0})}$ by (2)(2)( 2 ) as $\|M^{(\kappa_{0})}\|=\kappa_{0}$. Consequently, $P\cong\bigoplus\limits_{\lambda\in\Lambda}M^{(\kappa_{0})}\cong M^{(\kappa)}$.
Case II: $\|R\|+\aleph+\aleph_{0}<\kappa<\kappa_{0}.$ Let $P$ be an $R$-module with cardinal $\kappa$ in $\mathrm{Add}(M)$. Again, $P=\bigoplus\limits_{i\in I}P_{i}$ where each $P_{i}$ is $(\leq\kappa)$-generated $R$-module by Lemma 2.1. Since $\|P\|=\kappa>\|R\|+\aleph+\aleph_{0}$ and $\|P_{i}\|\leq\|R\|+\aleph+\aleph_{0}$ for each $i\in I$, we have $|I|=\kappa.$ Note that $\|M^{(\kappa)}\|=\kappa$. So $P\cong M^{(\kappa)}$ by (2).(2).( 2 ) . That is,
@@
Claim: There is a partition $\{I_{\lambda}\mid\lambda\in\Lambda\}$ (((resp., $\{J_{\lambda}\mid\lambda\in\Lambda\})$ of $I$ (((resp., $J)$ with $|I_{\lambda}|\leq\kappa$ (((resp., $|J_{\lambda}|\leq\kappa)$ for each $\lambda\in\Lambda$ such that
@@
If we set $M=R$, then $\mathrm{Add}(M)$ is the class of all projective modules. The above theorem recovers Trlifaj’s characterization of the categoricity for projective modules (see [9, Proposition 1.1]).
Recall that a class $\mathcal{D}$ of $R$-modules is called ($\kappa$-)deconstructible if there is a cardinal $\kappa$ such that every $R$-module $M$ in $\mathcal{D}$ is a transfinite extensions of modules from $\mathcal{C}$ consisting of $<\kappa$-presented modules, that is, $R$-modules $M$ that possess a $\mathcal{C}$-filtration, i.e., an increasing chain of submodules, $\mathcal{M}=(M_{\alpha}\mid\alpha\leq\sigma)$, such that $M_{0}=0$, $M_{\sigma}=M$, $M_{\alpha}=\bigcup_{\beta<\alpha}M_{\beta}$ for each limit ordinal $\alpha\leq\sigma$, and $M_{\alpha+1}/M_{\alpha}$ is isomorphic to an $R$-module in $\mathcal{C}$ for each $\alpha<\sigma$.
@@
Let $R$ be a ring, $S\SS$S the class of all semisimple $R$-modules and $S$ the direct sum of all simple modules non-isomorphic to each other. Then the following statements are equivalent:
@@
    $S\SS$S is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
@@
    $S\SS$S is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
@@
If we set $M$ to be the direct sum of all simple $R$-modules non-isomorphic to each other, then $S\mathrm{Add}(M)=\SS$S. So the equivalence of (1)(4)(1)-(4)( 1 ) - ( 4 ) follows by Theorem 2.2.
(5)(4)(5$\Rightarrow$(4)( 5 ) ⇒ ( 4 ) Suppose $R$ admits a unique simple module up to isomorphism. Then every semisimple $R$-module is $S$-free.
(1)(5)(1$\Rightarrow$(5)( 1 ) ⇒ ( 5 ) On contrary, there are two non-isomorphic simple $R$-modules $S_{1}$ and $S_{2}$. Then $S_{1}^{(\kappa)}\cong S_{2}^{(\kappa)}$ for all $\kappa>\|R\|+\aleph_{0}$, which is impossible. ∎
@@
    $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$.
@@
If we set $M$ to be the direct sum of all finitely presented $R$-modules non-isomorphic to each other, then $\mathrm{Add}(M)=P\mathcal{P}roj$. So the equivalence of (1)(4)(1)-(4)( 1 ) - ( 4 ) follows by Corollary 2.4.
(5)(1)(5$\Rightarrow$(1)( 5 ) ⇒ ( 1 ) Suppose $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$. Then every $R$-module is free, and so the result follows by Corollary 2.4.
(1)(5)(1$\Rightarrow$(5)( 1 ) ⇒ ( 5 ) Suppose $P\mathcal{P}roj$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$. Let $P$ be any nonzero pure-projective $R$-module. Then $P^{((\|R\|+\aleph_{0})^{+})}$ is pure-projective with $\|P^{((\|R\|+\aleph_{0})^{+})}\|\geq(\|R\|+\aleph_{0})^{+}$ and thus free by Corollary 2.4. Hence $P$ is a projective $R$-module. So every $R$-module is flat, that is, $R$ is a von Neumann regular ring. We claim that every idempotent in $R$ is trivial. Indeed, let $\{e_{\lambda}\mid\lambda\in\Lambda\}$ be a family of nonzero orthogonal idempotents in $R$. Let $\kappa>\|R\|+\aleph_{0}$. Then $(e_{\lambda}R)^{(\kappa)}$ is a pure-projective $R$-module whose cardinal is $\kappa$. Hence $(e_{\lambda}R)^{(\kappa)}\cong R^{(\kappa)}$. Consequently, $e_{\lambda}=1$. It follows by [2, Corollary 2.16] that $R$ is a semisimple ring. Consequently, $R\cong M_{n_{1}}(D_{1})\times\cdots\times M_{n_{k}}(D_{k})$ where each $D_{k}$ is a division ring and each $n_{k}\geq 1$. Since different $M_{n_{i}}(D_{i})^{(\kappa)}$s as semisimple $R$-modules have the same cardinal $\kappa$ but not isomorphic to each other, so $k=1$. Hence we have $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$. ∎

### v6

@@ Abstract
As an application, we show that the class of pure-projective $R$-modules is categorical in some (all) big cardinal if and only if the module $P^{(\aleph_{0})}$ is free for each countably generated pure-projective $R$-module $P$; the class of semisimple $R$-modules is categorical in some (all) big cardinal if and only if $R$ admits a unique simple module up to isomorphism, partly answering a question proposed in [5, Mazari-Armida M., Characterizing categoricity in several classes of modules. J. Algebra 617, 382-401 (2023)].
Key Words: $\mathrm{Add}(M)$, pure-projecitve modules, categoricity, Shelah’s Categoricity Conjecture.
@@ 1. Introduction
In this paper, $R$ is always an associative ring with unity and all modules are right modules unless otherwise stated. Let $S$ be a set, we write $|S|$ for its cardinal. Given an $R$-module $M$, we write $|M|$ for its underlying set and $\|M\|$ for its cardinal.
@@
((=Theorem 2.2)) Let $R$ be a ring and $M$ be an $R$-module which can be generated by $\leq\aleph$ elements. Then the following statements are equivalent:
@@
((=Theorem 2.5)) Let $R$ be a ring, $S\SS$ the class of all semisimple $R$-modules and $S$ the direct sum of all simple modules non-isomorphic to each other. Then the following statements are equivalent:
@@
    $S\SS$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
@@
    $S\SS$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
@@
((=Theorem 2.7)) Let $R$ be a ring and $P\mathcal{P}roj$ the class of all pure-projective $R$-modules. Then the following statements are equivalent:
@@
    $R$ is a von Neumann regular ring over which every projective $R$-module of cardinal $\kappa$ is free for all ((some)) $\kappa>\|R\|+\aleph_{0}$.
  6. (6)
    The module $P^{(\aleph_{0})}$ is free, for each countably generated pure-projective $R$-module $P$.
@@
Let $M$ be an $R$-module. Set $\mathrm{Add}(M)$ to be the class of all $R$-modules which is a direct summand of direct sums of copies of $M$. We call an $R$-module NN is $M$-free if $N\cong M^{(\kappa)}$ for some cardinal $\kappa.$ The following main theorem shows that the categoricity of $\mathrm{Add}{M}$ in a big cardinal is equivalent to its categoricity in a tail of cardinals.
@@
$(3)\Rightarrow(1)\Rightarrow(2)$ and $(3)\Rightarrow(4)\Rightarrow(2)$: Trivial.
$(2)\Rightarrow(3)$ Suppose $\mathrm{Add}(M)$ is $\kappa_{0}$-categorical in some fixed cardinal $\kappa_{0}>\|R\|+\aleph+\aleph_{0}$. We will show every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is $M$-free by dividing into two cases:
Case I: $\kappa\geq\kappa_{0}.$ Let $P$ be an $R$-module with cardinal $\kappa$ in $\mathrm{Add}(M)$. Then $P=\bigoplus\limits_{i\in I}P_{i}$ where each $P_{i}$ is $(\leq\kappa)$-generated $R$-module by Lemma 2.1. Since $\|P\|=\kappa>\|R\|+\aleph+\aleph_{0}$ and $\|P_{i}\|\leq\|R\|+\aleph+\aleph_{0}$ for each $i\in I$, we have $|I|=\kappa.$ Then we can give a partition $\{I_{\lambda}\mid\lambda\in\Lambda\}$ of $I$ with each $|I_{\lambda}|=\kappa_{0}$. Set $\bigoplus\limits_{i\in I_{\lambda}}P_{i}=P^{\prime}_{\lambda}$. Then $P=\bigoplus\limits_{\lambda\in\Lambda}P^{\prime}_{\lambda}$. Note that $\|P^{\prime}_{\lambda}\|=\kappa_{0}$. So $P^{\prime}_{\lambda}\cong M^{(\kappa_{0})}$ by $(2)$ as $\|M^{(\kappa_{0})}\|=\kappa_{0}$. Consequently, $P\cong\bigoplus\limits_{\lambda\in\Lambda}M^{(\kappa_{0})}\cong M^{(\kappa)}$.
Case II: $\|R\|+\aleph+\aleph_{0}<\kappa<\kappa_{0}.$ Let $P$ be an $R$-module with cardinal $\kappa$ in $\mathrm{Add}(M)$. Again, $P=\bigoplus\limits_{i\in I}P_{i}$ where each $P_{i}$ is $(\leq\kappa)$-generated $R$-module by Lemma 2.1. Since $\|P\|=\kappa>\|R\|+\aleph+\aleph_{0}$ and $\|P_{i}\|\leq\|R\|+\aleph+\aleph_{0}$ for each $i\in I$, we have $|I|=\kappa.$ Note that $\|M^{(\kappa)}\|=\kappa$. So $P\cong M^{(\kappa)}$ by $(2).$ That is,
@@
Claim: There is a partition $\{I_{\lambda}\mid\lambda\in\Lambda\}$ ((resp., $\{J_{\lambda}\mid\lambda\in\Lambda\})$ of II ((resp., $J)$ with $|I_{\lambda}|\leq\kappa$ ((resp., $|J_{\lambda}|\leq\kappa)$ for each $\lambda\in\Lambda$ such that
@@
If we set $M=R$, then $\mathrm{Add}(M)$ is the class of all projective modules. The above theorem recovers part of Trlifaj’s characterization of the categoricity for projective modules (see [9, Proposition 1.1]).
Recall that a class $\mathcal{D}$ of $R$-modules is called $(\kappa$-))deconstructible if there is a cardinal $\kappa$ such that every $R$-module $M$ in $\mathcal{D}$ is a transfinite extensions of modules from $\mathcal{C}$ consisting of $\kappa-$presented modules, that is, $R$-modules $M$ that possess a $\mathcal{C}$-filtration, i.e., an increasing chain of submodules, $\mathcal{M}=(M_{\alpha}\mid\alpha\leq\sigma)$, such that $M_{0}=0$, $M_{\sigma}=M$, $M_{\alpha}=\bigcup_{\beta<\alpha}M_{\beta}$ for each limit ordinal $\alpha\leq\sigma$, and $M_{\alpha+1}/M_{\alpha}$ is isomorphic to an $R$-module in $\mathcal{C}$ for each $\alpha<\sigma$.
@@
Let $R$ be a ring, $S\SS$ the class of all semisimple $R$-modules and $S$ the direct sum of all simple modules non-isomorphic to each other. Then the following statements are equivalent:
@@
    $S\SS$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
@@
    $S\SS$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
@@
[…]
$(5)\Rightarrow(4)$ Suppose $R$ admits a unique simple module up to isomorphism. Then every semisimple $R$-module is $S$-free.
$(1)\Rightarrow(5)$ On contrary, there are two non-isomorphic simple $R$-modules $S_{1}$ and $S_{2}$. Then $S_{1}^{(\kappa)}\cong S_{2}^{(\kappa)}$ for all $\kappa>\|R\|+\aleph_{0}$, which is impossible. ∎
@@
    $R$ is a von Neumann regular ring over which every projective $R$-module of cardinal $\kappa$ is free for all ((some)) $\kappa>\|R\|+\aleph_{0}$.
  6. (6)
    The module $P^{(\aleph_{0})}$ is free, for each countably generated pure-projective $R$-module $P$.
@@
If we set $M$ to be the direct sum of all finitely presented $R$-modules non-isomorphic to each other, then $\mathrm{Add}(M)=P\mathcal{P}roj$. So the equivalence of (1)(4)(1)-(4) follows by Corollary 2.4.
$(5)\Rightarrow(1)$ Suppose $R$ is a von Neumann regular ring. Then every pure-projective $R$-module is projective. And so the result follows by Corollary 2.4.
$(1)\Rightarrow(5)$ Suppose $P\mathcal{P}roj$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$. Let $P$ be any nonzero pure-projective $R$-module. Then $P^{((\|R\|+\aleph_{0})^{+})}$ is pure-projective with $\|P^{((\|R\|+\aleph_{0})^{+})}\|\geq(\|R\|+\aleph_{0})^{+}$ and thus free by Corollary 2.4. Hence $P$ is a projective $R$-module. So every $R$-module is flat, that is, $R$ is a von Neumann regular ring. It follows by Corollary 2.4 that every pure-projective $R$-module of cardinal $\kappa$ is free for all $\kappa>\|R\|+\aleph_{0}$. Since every pure-projective module is projective over von Neumann regular rings. Hence, every projective $R$-module of cardinal $\kappa$ is free for all (some) $\kappa>\|R\|+\aleph_{0}$.
$(2)\Rightarrow(6)$ Let $P$ be a nonzero countably generated pure-projective $R$-module. Since $\kappa>\|R\|+\aleph_{0}$, both $R^{(\kappa)}$ and $P^{(\kappa)}$ have cardinality $\kappa$. So we have an isomorphism $P^{(\kappa)}\cong R^{(\kappa)}$, and hence $P$ is projective. Consequently, the countably generated $R$-module $P^{(\aleph_{0})}$ is isomorphic to a direct summand of $R^{(\aleph_{0})}$, i.e., $P^{(\aleph_{0})}\oplus M\cong R^{(\aleph_{0})}$ for an $R$-module $M$. Similarly, $R^{(\aleph_{0})}\oplus N\cong P^{(\aleph_{0})}$ for an $R$-module NN. Then
$P^{(\aleph_{0})}\cong(R^{(\aleph_{0})}\oplus R^{(\aleph_{0})})\oplus N\cong R^{(\aleph_{0})}\oplus P^{(\aleph_{0})}$
and similarly $R^{(\aleph_{0})}\oplus P^{(\aleph_{0})}\cong R^{(\aleph_{0})}$. Consequently, $P^{(\aleph_{0})}\cong R^{(\aleph_{0})}$ is free.
$(6)\Rightarrow(4)$ Let $P$ be pure-projective module of cardinal $(\|R\|+\aleph_{0})^{+}$. Then $P$ is a directed sum of $(\|R\|+\aleph_{0})^{+}$ copies of countably generated pure-projective $R$-modules. It follows by $(6)$ that $P$ is free. ∎